OpenStudy (anonymous):

find all values of x at which the tangent line is horizontal at y=(x^2-1)/(x+2)

OpenStudy (across):

Dannibe7, do you know what a horizontal tangent line is?

OpenStudy (anonymous):

where the slope of the tan line is zero?

OpenStudy (across):

I presume you know how to differentiate?

OpenStudy (anonymous):

i got that the derivative is (x^2+4x+1)/(x+2)^2 but what do i do to find where it is horizontal?

OpenStudy (across):

By definition, the derivative of a function is the slope of the tangent line at any point on this function. When is a slope horizontal?

OpenStudy (anonymous):

when it is zero. but i dont know how to get it to that point after i find the derivative

OpenStudy (across):

That's correct!\[f(x)=\frac{x^2-1}{x+2}\]\[f'(x)=\frac{x^2+4x+1}{(x+2)^2}\]Now, since we know the slope has to be zero, all we have to do is solve the following equation:\[\frac{x^2+4x+1}{(x+2)^2}=0.\]Am I correct?

OpenStudy (anonymous):

thats what I thought to do but i cant come up with the right answer the answer is supposed to be \[-2\pm \sqrt{3}\]

OpenStudy (anonymous):

correct ans

OpenStudy (across):

\[\frac{x^2+4x+1}{(x+2)^2}=0,\]\[x^2+4x=-1,\]\[x^2+4x+4=-1+4,\]in that step above, I simply completed the squares,\[(x+2)^2=3,\]\[x+2=\pm\sqrt{3}\]\[x=-2\pm\sqrt{3}.\]All I did was algebra.

OpenStudy (anonymous):

what is completing the square?

OpenStudy (across):

Since there are sources that can do a better job than me at explaining what it is, I will link you to this website: http://www.purplemath.com/modules/sqrquad.htm